# Chance-Constrained DC Optimal Power Flow Using Constraint-Informed Statistical Estimation

### v1

@@ IV-C Out-of-Sample Risk and System Reliability
To evaluate the quality of solutions obtained, we use 20% of the dataset as a holdout, and denote it as $\mathcal{D}$. The empirical violation of a single chance constraint (i.e., generation or line flow limit) is then defined as [12]:
@@
In the Synthetic-G dataset, the positive log-likelihood in Table III indicates that the obtained Gaussian parameters are a good fit to the forecast errors. This directly translates to the constraint violations being below the targeted $\epsilon=0.05$.
In the Synthetic-C dataset, we observe that both approaches yield worst-case constraint violations that exceed $\epsilon$. The violations in the classical approach average roughly $0.5$, whereas the constraint-informed approach achieves much smaller violations averaging around $0.1$ with lower overall variance across the ten runs. Similar patterns can be seen in the NordPool dataset, where constraint-informed estimation consistently results in lower violations and variability compared to the classical approach.
We highlight that both approaches can sometimes produce statistical parameters that lead to infeasible optimization models, as shown in Figure 5. The constraint-informed estimation, however, results in fewer infeasible models. We find that the infeasibility occurs because the estimated parameters violate constraints (33) and (36). This infeasibility can be resolved using a modified FixedEM algorithm [36] to constrain the means of the GMM components at zero (see Appendix C).
Figure 4: Out-of-sample worst-case constraint violations.
Figure 5: Number of infeasible optimization models (out of ten) in the Synthetic-C and NordPool datasets.
@@ IV-D Computational Time
We compare total computation times for GMM estimation ($K=3$) on the NordPool dataset. Table IV reports the total time across 10 runs of EM, each with a different random initialization (and under both spherical and tied covariances in case of $\bm{\eta}_{l}$), and then selecting the best model using the Bayesian Information Criterian (BIC) score. We find that the computational times are small and suitable for practical real-time deployment. Unsurprisingly, the optimization time is identical for both approaches.
TABLE IV: Computational time comparisons of the two approaches
@@
Statistical Fitting 0.04s for $\bm{\xi}$ 0.05s for $\Omega$ and 0.09s per $\bm{\eta}_{l}$
Optimization 0.03s 0.03s
@@ V Conclusions
We study the interplay between statistics and optimization in the chance-constrained optimal power flow (CC-OPF) problem under uncertainties arising from wind power forecast errors. Our proposed constrained-informed approach uses the OPF constraints to isolate the true relevant uncertainties to two sources of randomness: a one-dimensional aggregate system-wide forecast error and a two-dimensional line-specific flow error. By reducing dimensionality, our approach integrates statistical estimation with chance-constrained optimization, thus providing system operators with information to make more reliable dispatch decisions.
@@ Appendix A Equivalence of Classical and Constraint-Informed Gaussian MLE
$\hat{\bm{\mu}}=\frac{1}{N}\sum_{j=1}^{N}\bm{x}^{(j)},\;\;\hat{\bm{\Sigma}}=\frac{1}{N}\sum_{j=1}^{N}\bigl{(}\bm{x}^{(j)}-\hat{\bm{\mu}}\bigr{)}\bigl{(}\bm{x}^{(j)}-\hat{\bm{\mu}}\bigr{)}^{\top}$. (40)
In the constraint-informed approach, we instead transform the data first to obtain the lower‐dimensional samples:
@@ Appendix B Constraint-Informed CC-OPF Reformulation
Appendix C Effects of Zeroing GMM Component Means
In this section, we use a modified FixedEM algorithm [36] to explicitly constrain $\hat{\bm{\mu}}_{k}=\bm{0}$ across all Gaussian components. As seen in Table V, this benefits the classical approach in the Synthetic-C dataset by substantially reducing the log-likelihood gap to the constraint-informed approach (compare to Table III). The latter sees almost no improvement, yet it still outperforms the classical fit. For the NordPool dataset, whose distribution is not centered near zero, zeroing the means reduces the performance of the constraint-informed fit; in contrast, zeroing means improves the classical fit since this reduces the search space for EM by ten dimensions in each component. Indeed, in both datasets, fixing the means improves performance of the (non-global) EM algorithm.
TABLE V: Log-likelihood of $\Omega$ with zero means GMM
Dataset, Fitted Model Classical Constraint-Informed
Synthetic-C, GMM ($K=3$)
-9978 -9868
NordPool, GMM ($K=3$)
-3584 -3403
The improvement in estimation also helps resolve infeasibility in optimization. Specifically, only one and zero runs in the classical and constraint-informed cases, respectively, are infeasible. Also, the average worst-case constraint violation in the classical case improves from roughly $0.5$ to $0.2$, for both Synthetic-C and NordPool datasets, although we emphasize that this is still worse than the constraint-informed approach. For the latter, zeroing the means is not a good choice given the skewed nature of the NordPool dataset, and it also leads to a slightly worse out-of-sample risk. Nonetheless, for the classical approach, the potential misfit from fixing the means at zero still improves the quality of solutions compared to those obtained using the traditional EM algorithm.

### v2

@@ IV-C Out-of-Sample Risk and System Reliability
As noted in Section III-E, unconstrained EM may produce component mean estimates that violate (34) and (37), rendering the optimization model infeasible. To prevent this, all optimization-related results in this and subsequent sections use the FixedEM algorithm [36], which constrains every component mean to zero for both the classical ($\bm{\xi}$) and the constraint-informed approach ($\Omega,\bm{\eta}_{l}$). For each $\bm{\eta}_{l}$ fitting, we fix $K=3$ and fit under both tied and spherical covariance structures with 10 initializations each; because the zero-mean constraint reduces the effective model complexity, BIC does not favor additional components beyond $K=3$.
To evaluate the quality of CC-OPF solutions obtained, we use 20% of the dataset as a holdout, and denote it as $\mathcal{D}$. The empirical violation of a single chance constraint (i.e., generation or line flow limit) is then defined as [12]:
@@
In the Synthetic-G dataset, the negative BIC score in Table III indicates that the obtained Gaussian parameters are a good fit to the forecast errors. This directly translates to the constraint violations being below the targeted $\epsilon=0.05$.
In the Synthetic-C dataset, we observe that both approaches yield worst-case constraint violations that exceed $\epsilon$. The violations in the classical approach average roughly $0.2$, whereas the constraint-informed approach achieves much smaller violations, averaging slightly below $0.1$ with lower overall variance across the ten runs.
Similar patterns can be seen in the NordPool dataset, where constraint-informed estimation consistently results in lower violations and variability compared to the classical approach. The NordPool dataset exhibits both skewness and a limited number of samples, leading to greater variability across runs even under the constraint-informed approach. We also note that constraint-informed violations are higher on NordPool than on Synthetic-C, suggesting that enforcing zero-mean components may be less effective when the underlying distribution is significantly skewed.
Overall, these results demonstrate that the constraint-informed dimensionality reduction consistently improves out-of-sample reliability: by fitting GMMs directly to the low-dimensional variables that enter each chance constraint, the EM algorithm better captures heavy-tailed (Synthetic-C) and skewed (NordPool) forecast errors, translating the improved goodness-of-fit into lower and less variable constraint violations.
Figure 4: Out-of-sample worst-case constraint violations across ten runs. The right panel corresponds to the test system discussed in Section IV-E, where the classical approach produced infeasible optimization models in all runs.
@@ IV-D Computational Time
Table IV reports GMM fitting times on the NordPool dataset using the FixedEM algorithm. For $\bm{\xi}$ and $\Omega$, we sweep $K\in\{1,\dots,10\}$ with 10 random initializations per $K$ and select the best model by BIC. For each $\bm{\eta}_{l}$, we fix $K=3$ and fit under both tied and spherical covariance structures with 10 initializations each. The per-line fitting time is consequently much smaller than that of $\Omega$, since fewer candidate models are evaluated. Due to the small size of the network, the optimization time for both approaches is relatively small. For the constraint-informed approach, the total computational time is 19.0 seconds, which is suitable for practical real-time deployment. Moreover, the per-line fits for $\bm{\eta}_{l}$ are independent and can be trivially parallelized.
TABLE IV: Computational time comparisons on the IEEE 118-Bus System
@@
Statistical Fitting 2.81s for $\bm{\xi}$ 2.61s for $\Omega$ and 0.09s per $\bm{\eta}_{l}$
Optimization 0.34s 0.43s
IV-E Scalability
To evaluate the scalability of the constraint-informed approach, we consider the 2746-bus Polish 2003–2004 winter peak case [40]. Following [6], we simulate 20% wind penetration (of total load), evenly distributed across 50 wind units co-located with the 50 largest controllable generators. We use the Synthetic-C dataset where each wind unit is assigned Cauchy-distributed forecast errors. [^6] Cauchy parameters: $x_{0}=0,\gamma=0.01$. The scale parameter $\gamma$ is reduced from $0.02$ in the IEEE-118 case to moderate the forecast error magnitude for this larger network.
Figure 5 plots the best BIC score as a function of the number of Gaussian components $K$ for both approaches as the uncertainty dimension increases from 10 to 50. As in Section IV-B, all goodness-of-fit comparisons in Figure 5 use classical EM to isolate the effect of dimensionality on statistical accuracy. The constraint-informed approach consistently achieves lower (better) BIC values across all $K$, and its 50-dimensional curve nearly overlaps with the classical fit at 10 dimensions. Moreover, the constraint-informed BIC stabilizes around $K=3$ or $4$ even at high dimensions, indicating that a small number of components suffices to capture the distribution of $\Omega$. In contrast, the classical approach does not exhibit a clear BIC minimum within the examined range; at $K\leq 2$ in 50 dimensions, the fitted GMM assigns negligible density to tail observations, causing a collapse in the log-likelihood and a sharp increase in BIC.
Figure 5: Best BIC score versus $K$ for the aggregate forecast error $\Omega$ at different uncertainty dimensions on the Polish 2746-bus system.
Table V reports fitting and optimization times on the Polish system using the FixedEM algorithm, following the same procedure as in Section IV-D. The optimization times are averaged across all 10 runs irrespective of infeasibility status. As the uncertainty dimension grows from 10 to 50, the classical fitting time increases substantially, whereas the constraint-informed fitting time for $\Omega$ remains nearly unchanged since it always performs univariate estimation. The main computational cost of the constraint-informed approach lies in the per-line fits for $\bm{\eta}_{l}$; however, the per-line fitting time is roughly constant regardless of network size, and these fits are independent and can be fully parallelized. Even when performed sequentially, the total fitting and optimization time is 531.0 seconds, which is still within a typical OPF dispatch window. On the optimization side, the classical approach yields infeasible models due to poor high-dimensional parameter estimates, while the constraint-informed approach produces a feasible solution within minutes. These results confirm that the constraint-informed approach scales to large systems, maintaining both statistical accuracy and computational tractability.
TABLE V: Computational time comparisons on the Polish 2746-bus system
Classical Constraint-Informed
Statistical Fitting 12.27s for $\bm{\xi}$ 3.01s for $\Omega$ and 0.13s per $\bm{\eta}_{l}$
Optimization 183.97s 101.68s
The out-of-sample worst-case constraint violation for the constraint-informed approach is also illustrated in Figure 4. The stability of the low-dimensional fitting indicates that the worst-case violation remains largely unchanged as the network scales from 118 buses to 2746 buses. In contrast, the classical approach yields infeasible models across all ten runs, due to estimation challenges in fitting 50-dimensional GMMs.
@@ V Conclusions
We study the interplay between statistics and optimization in the chance-constrained optimal power flow (CC-OPF) problem under uncertainties arising from wind power forecast errors. Our proposed constraint-informed approach uses the OPF constraints to isolate the true relevant uncertainties to two sources of randomness: a one-dimensional aggregate system-wide forecast error and a two-dimensional line-specific flow error. By reducing dimensionality, our approach integrates statistical estimation with chance-constrained optimization, thus providing system operators with information to make more reliable dispatch decisions.
@@
A limitation of the current approach is that it fits separate low-dimensional distributions for each constraint-relevant projection rather than constructing a single globally coherent low-dimensional model. While this does not affect the enforcement of individual (or even two-sided) chance constraints, it would need to be addressed in settings that require joint chance constraints coupling multiple generators or lines. Extending the constraint-informed framework to such joint formulations, as well as to time-coupled problems such as unit commitment, is a promising direction for future work.
@@ Appendix A Equivalence of Classical and Constraint-Informed Gaussian MLE
[…]
In the constraint-informed approach, we instead transform the data first to obtain the lower-dimensional samples: